# Unscented Kalman Filter with a Nonlinear Propagation Model for Navigation Applications

### v3

@@ III Proposed Method
where $\oplus$ denotes addition via transformation matrix and ${x_{\sigma}}_{k}$ is a representative solution of the sigma point $\sigma$. We shall compute ${x_{\sigma}}_{k+1}$ using the navigation algorithm cycle ((24) - (32)) and label it na_cyclena\_cycle. To apply the navigation algorithm cycle, from step $k$ to step $k+1$, the accelerations and gyroscopes measurements $f_{ib}^{b},w_{ib}^{b}$ (For brevity, we omit the subscript $ib$ in the rest of the derivation) are required. Since the sigma points hold also the estimated bias errors, we need to compensate the measurements with the appropriate biases before applying the navigation cycle, reflecting the sigma point states. Thus using (35) - (37) we can write:
@@ IV-B Evaluation metrics and baseline
where, $n$ is the number of samples in the trajectory, $\|{\delta\boldsymbol{v}(j)}\|$ is the norm of the velocity error between the estimated velocity and the GT in time step $i$. The track average is:
@@
where $VRMSE_{5}$ and $VRMSE_{6}$ are the $VRMSE$ of tracks 5 and 6, respectively.
@@
where $\delta{\boldsymbol{\Psi}}(j)$} is the misalignment vector that represents by the Euler angles corresponding to the following transformation matrix $\boldsymbol{\mathbf{C}}_{n_{r}}^{n}$:

### v4

@@ III Proposed Method
where $\oplus$ denotes addition via transformation matrix and ${x_{\sigma}}_{k}$ is a representative solution of the sigma point $\sigma$. We shall compute ${x_{\sigma}}_{k+1}$ using the navigation algorithm cycle ((24) - (32)) and label it na_cyclena\_cycle. To apply the navigation algorithm cycle, from step $k$ to step $k+1$, the accelerations and gyroscopes measurements $f_{ib}^{b},w_{ib}^{b}$ (For brevity, we omit the subscript ibib in the rest of the derivation) are required. Since the sigma points hold also the estimated bias errors, we need to compensate the measurements with the appropriate biases before applying the navigation cycle, reflecting the sigma point states. Thus using (35) - (37) we can write:
@@ IV-B Evaluation metrics and baseline
where, $n$ is the number of samples in the trajectory, $\|{\delta\boldsymbol{v}(i)}\|$ is the norm of the velocity error between the estimated velocity and the GT in time step ii. The track average is:
@@
where $VRMSE_{5}$ and $VRMSE_{6}$ are the VRMSEVRMSE of tracks 5 and 6, respectively.
@@
where $\delta{\boldsymbol{\Psi}}(i)$} is the misalignment vector that represents by the Euler angles corresponding to the following transformation matrix $\boldsymbol{\mathbf{C}}_{n_{r}}^{n}$: